# Autocratic strategies in Cournot oligopoly game

### v2

@@ Abstract
An oligopoly is a market in which the price of a goods is controlled by a few firms. Cournot introduced the simplest game-theoretic model of oligopoly, where profit-maximizing behavior of each firm results in market failure. Furthermore, when the Cournot oligopoly game is infinitely repeated, firms can tacitly collude to monopolize the market. Such tacit collusion is realized by the same mechanism as direct reciprocity in the repeated prisoner’s dilemma game, where mutual cooperation can be realized whereas defection is favorable for both prisoners in one-shot game. Recently, in the repeated prisoner’s dilemma game, a class of strategies called zero-determinant strategies attracts much attention in the context of direct reciprocity. Zero-determinant strategies are autocratic strategies which unilaterally control payoffs of players. There were many attempts to find zero-determinant strategies in other games and to extend them so as to apply them to broader situations. In this paper, first, we show that zero-determinant strategies exist even in the repeated Cournot oligopoly game. Especially, we prove that an averagely unbeatable zero-determinant strategy exists, which is guaranteed to obtain the average payoff of the opponents. Second, we numerically show that the averagely unbeatable zero-determinant strategy can be used to promote collusion when it is used against an adaptively learning player, whereas it cannot promote collusion when it is used against two adaptively learning players. Our findings elucidate some negative impact of zero-determinant strategies in oligopoly market.
@@
Repeated games have been used to analyze the rational decision-making of multiple agents in a long-term interdependent relationship. Recently, a class of autocratic strategies, called zero-determinant strategies, was discovered in repeated games, which unilaterally controls payoffs of players. So far, properties of zero-determinant strategies in social dilemma situations have extensively been investigated, and it has been shown that some zero-determinant strategies promote cooperation. Moreover, zero-determinant strategies have been found in several games. However, it has not been known whether zero-determinant strategies exist in oligopoly games. In this paper, we investigate zero-determinant strategies in the repeated Cournot oligopoly game, which is the simplest mathematical model of oligopoly. We prove the existence of zero-determinant strategies which unilaterally enforce linear relations between the payoff of the player and the average payoff of the opponents with both positive and negative slopes. Furthermore, we numerically show that some zero-determinant strategy can promote collusion in duopoly case, although it cannot promote collusion in triopoly case. Our results imply that zero-determinant strategies can be used to promote cooperation between firms even in oligopoly market.
@@ 1 Introduction
Oligopoly is one of the simplest game-theoretic situations in economics. Cournot introduced a simple model of oligopoly, where multiple firms produce the same goods and its price decreases as the total amount of the goods increases [1, 2]. Each firm wants to maximize its profit, but needs to choose its production while considering production of the other firms. When all firms are enough rational, the Cournot-Nash equilibrium is realized, where production of each firm is the best response against production of the other firms. Because total production in this equilibrium is smaller than that in perfect competition, oligopoly results in market failure.
@@
When the Cournot oligopoly game is infinitely repeated, situation becomes worse for consumers. Since firms become taking total profits obtained in future into account, firms can tacitly collude to monopolize the market [3], which is an example of the folk theorem in repeated games. This result is similar to direct reciprocity in the repeated prisoner’s dilemma game [4]. In the prisoner’s dilemma game, two prisoners independently choose cooperation or defection. Defection is the best action in terms of the payoff of each prisoner regardless of the action of the other prisoner, and mutual defection is realized as a result of rational behavior. Because mutual cooperation improves the payoffs of both prisoners in mutual defection, this game describes a kind of social dilemma. However, when the game is infinitely repeated, mutual cooperation can be realized as a result of long-term perspectives, similarly to tacit collusion in the repeated Cournot oligopoly game.
@@
In 2012, a novel class of strategies, called zero-determinant (ZD) strategies, was discovered in the repeated prisoner’s dilemma game [5]. Counterintuitively, ZD strategies unilaterally control payoffs of players by enforcing linear relations between payoffs. ZD strategies can be regarded as generalization of the equalizer strategy [6], which unilaterally sets the payoff of the opponent. ZD strategies also contain the extortionate ZD strategy, by which a player never obtains payoff lower than that of the opponent, and the generous ZD strategy [7], by which a player never obtains payoff higher than that of the opponent but promotes cooperation. So far, ZD strategies were discovered in the prisoner’s dilemma game [5, 7, 8], the public goods game [9, 10], the continuous donation game [11], the asymmetric prisoner’s dilemma game [12, 13], two-player potential games [14, 15], and symmetric games with no generalized rock-paper-scissors cycles [16, 17]. Particularly, ZD strategies in games with infinite action sets are called as autocratic strategies [11]. Furthermore, ZD strategies were extended to broader situations such as games with a discount factor [18, 11, 19, 20], games with observation errors [21, 22, 23], asynchronous games [24], and stochastic games [25, 26]. Roles of ZD strategies in the context of evolution of cooperation has been extensively investigated [27, 28, 7, 29, 30, 31, 8, 32, 33].
@@
Although the Cournot oligopoly game contains some social dilemma structure as noted above, it is not completely the same as social dilemma. This is because the action sets in the Cournot oligopoly game are infinite sets, whereas the action sets in typical social dilemma games contain only two actions, that is, cooperation and defection. Probably for that reason, ZD strategies have not been discovered in the Cournot oligopoly game, and it is not clear whether players can autocratically control payoffs in the game.
@@
In this paper, we make two key contributions. First, we prove that ZD strategies (or, autocratic strategies) also exist in the repeated Cournot oligopoly game. Concretely, we show that ZD strategies pinning their own payoffs, positively correlated ZD strategies [34, 35], which are the class containing the extortionate ZD strategy and the generous ZD strategy, and negatively correlated ZD strategies exist in the game. Especially, we find that an averagely unbeatable ZD strategy exists in the game, by which a player unilaterally obtains the average payoff of the opponents. Second, we specify performance of the averagely unbeatable ZD strategy in evolution of collusion. We numerically investigate whether collusion is achieved when the averagely unbeatable ZD strategy is used against adaptively learning players [5, 34, 35]. We find that, in a two-player case, the averagely unbeatable ZD strategy promotes collusion of a learning player, but in a three-player case, it cannot promote collusion.
@@
This paper is organized as follows. In Section 2, we introduce the Cournot oligopoly game and its repeated version. In Section 3, we introduce autocratic strategies, and explain their properties related to the existence. In Section 4, we prove that some types of autocratic strategies exist in the repeated Cournot oligopoly game. In Section 5, we provide numerical results. Section 6 is devoted to discussion.
@@ 2 Model
We consider the $N$-player Cournot oligopoly game with $N\geq 2$ [1, 2]. The action space of player $j$ is given by $A_{j}:=[0,x_{\mathrm{max}}]$ with $0<x_{\mathrm{max}}<\infty$ for all $j$. The action of player $j$ is described as $x_{j}\in A_{j}$, which represents production of the goods by player $j$. We collectively write the action profile as $\bm{x}:=\left(x_{1},\cdots,x_{N}\right)$. We also use the notation $\mathcal{A}:=\prod_{j}A_{j}$, $A_{-j}:=\prod_{k\neq j}A_{k}$, and $x_{-j}:=\left(x_{1},\cdots,x_{j-1},x_{j+1},\cdots,x_{N}\right)\in A_{-j}$. When we want to emphasize $x_{j}$ in $\bm{x}$, we write $\bm{x}=(x_{j},x_{-j})$. The payoff of player $j$ when the action profile is $\bm{x}$ is described as
@@
where $a>c>0$, $b>0$, and $\theta(\cdot)$ is the step function. That is, $\left[a-b\sum_{k=1}^{N}x_{k}\right]\theta\left(a-b\sum_{k=1}^{N}x_{k}\right)$ represents a price of the goods, and $c$ represents a cost of production. The price is a decreasing function of the total production $\sum_{k=1}^{N}x_{k}$. We assume that $x_{\mathrm{max}}\geq a/b$.
It is known that the Nash equilibrium of this game is $x_{j}=x^{(\mathrm{N})}:=\frac{a-c}{(N+1)b}$ and $s_{j}\left(\bm{x}^{(\mathrm{N})}\right)=\frac{(a-c)^{2}}{(N+1)^{2}b}$ for all $j$. On the other hand, if all players collude to monopolize the market, and share the payoffs equally, the realized state is $x_{j}=x^{(\mathrm{C})}:=\frac{a-c}{2Nb}$ and $s_{j}\left(\bm{x}^{(\mathrm{C})}\right)=\frac{(a-c)^{2}}{4Nb}$ for all $j$. Because $s_{j}\left(\bm{x}^{(\mathrm{C})}\right)>s_{j}\left(\bm{x}^{(\mathrm{N})}\right)$, the game can be regarded as one example of social dilemma. In addition, the Walras equilibrium, which corresponds to perfect competition, is $x_{j}=x^{(\mathrm{W})}:=\frac{a-c}{Nb}$ and $s_{j}\left(\bm{x}^{(\mathrm{W})}\right)=0$, where the price is equal to the marginal cost.
We consider a repeated version of the Cournot oligopoly game. We write the probability measure of an action profile at time $t$ by $P_{t}\left(\cdot\right)$. Here we consider the case that there is no discounting of future payoffs $(\delta=1)$: that is, the payoff of player $j$ in the repeated game is given by
@@
where $\left\langle\cdots\right\rangle^{*}$ is the expected value of the quantity $\cdots$ with respect to the time-average probability of an action profile
$P^{*}\left(\cdot\right)$ $:=\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}P_{t}\left(\cdot\right).$ (3)
@@ 3 Preliminaries
A memory-one strategy of player $j$ is a autocratic strategy when its strategy $T_{j}$ can be written in the form
[…]
with some coefficients $\left\{\alpha_{b}\right\}$ and some bounded function $\psi(\cdot)$.
The autocratic strategy (4) unilaterally enforces a linear relation between payoffs
00 0 $=\sum_{k=0}^{N}\alpha_{k}\left\langle s_{k}\right\rangle^{*}.$ (5)
Therefore, autocratic strategies can be used to control payoffs unilaterally. In this paper, we use two words “ZD strategy” and “autocratic strategy” interchangeably.
@@
00 0 $\leq B\left(\overline{x}_{j},x_{-j}\right)\leq W\quad(\forall x_{-j}\in A_{-j}).$ (6)
@@
[…]
with $\psi\left(\underline{x}_{j}\right)=\psi_{0}$ and $\psi\left(\overline{x}_{j}\right)=\psi_{0}+W$. McAvoy and Hauert called such autocratic strategies two-point autocratic strategies, because such autocratic strategy uses only two actions.
Although they found that the condition (6) is a sufficient condition for the existence of autocratic strategies, Ueda proved that it is also a necessary condition when the number of actions is finite [16]. However, we do not know whether it is also a necessary condition when the number of actions is infinite. At this stage, we only prove the following proposition about two-point autocratic strategies.
@@
If a two-point autocratic strategy of player $j$ controlling $B$ exists, then there exist two actions $\underline{x}_{j},\overline{x}_{j}\in A_{j}$ such that

### v3

@@ Abstract
An oligopoly is a market in which the price of goods is controlled by a few firms. Cournot introduced the simplest game-theoretic model of oligopoly, where profit-maximizing behavior of each firm results in market failure. Furthermore, when the Cournot oligopoly game is infinitely repeated, firms can tacitly collude to monopolize the market. Such tacit collusion is realized by the same mechanism as direct reciprocity in the repeated prisoner’s dilemma game, where mutual cooperation can be realized whereas defection is favorable for both prisoners in a one-shot game. Recently, in the repeated prisoner’s dilemma game, a class of strategies called zero-determinant strategies attracts much attention in the context of direct reciprocity. Zero-determinant strategies are autocratic strategies which unilaterally control payoffs of players by enforcing linear relationships between payoffs. There were many attempts to find zero-determinant strategies in other games and to extend them so as to apply them to broader situations. In this paper, first, we show that zero-determinant strategies exist even in the repeated Cournot oligopoly game, and that they are quite different from those in the repeated prisoner’s dilemma game. Especially, we prove that a fair zero-determinant strategy exists, which is guaranteed to obtain the average payoff of the opponents. Second, we numerically show that the fair zero-determinant strategy can be used to promote collusion when it is used against an adaptively learning player, whereas it cannot promote collusion when it is used against two adaptively learning players. Our findings elucidate some negative impact of zero-determinant strategies in the oligopoly market.
@@
Repeated games have been used to analyze the rational decision-making of multiple agents in a long-term interdependent relationship. Recently, a class of autocratic strategies, called zero-determinant strategies, was discovered in repeated games, which unilaterally controls payoffs of players via enforcing linear relations between payoffs. So far, properties of zero-determinant strategies in social dilemma situations have extensively been investigated, and it has been shown that some zero-determinant strategies promote cooperation. Moreover, zero-determinant strategies have been found in several games. However, it has not been known whether zero-determinant strategies exist in oligopoly games. In this paper, we investigate zero-determinant strategies in the repeated Cournot oligopoly game, which is the simplest mathematical model of oligopoly. We prove the existence of zero-determinant strategies which unilaterally enforce linear relations between the payoff of the player and the average payoff of the opponents. Furthermore, we numerically show that some zero-determinant strategy can promote collusion in a duopoly case, although it cannot promote collusion in a triopoly case. Our results imply that zero-determinant strategies can be used to promote cooperation between firms even in an oligopoly market.
@@ 1 Introduction
Oligopoly is one of the simplest game-theoretic situations in economics, where the price of goods is controlled by a few firms. Cournot introduced a simple model of oligopoly, where multiple firms produce the same goods and the good’s price decreases as the total amount of the goods increases [1, 2]. Each firm wants to maximize its profit, but needs to choose its production while considering production of the other firms. When all firms are rational, the Cournot-Nash equilibrium is realized, where production of each firm is the best response against production of the other firms. Because total production in this equilibrium is smaller than that in perfect competition, oligopoly results in market failure.
@@
When the Cournot oligopoly game is infinitely repeated, the situation becomes worse for consumers. Since firms become taking total profits obtained in future into account, firms can tacitly collude to monopolize the market [3], even if the Cournot-Nash equilibrium production is the only rational behavior of each firm in a one-shot game. This is an example of the folk theorem in repeated games. This result is similar to direct reciprocity in the repeated prisoner’s dilemma game [4]. In the prisoner’s dilemma game, two prisoners independently choose cooperation or defection. Defection is the best action in terms of the payoff of each prisoner regardless of the action of the other prisoner, and mutual defection is realized as a result of rational behavior. Because mutual cooperation improves the payoffs of both prisoners in mutual defection, this game describes a kind of social dilemmas. However, when the game is infinitely repeated, mutual cooperation can be realized as a result of long-term perspectives, similarly to tacit collusion in the repeated Cournot oligopoly game.
@@
In 2012, a novel class of strategies, called zero-determinant (ZD) strategies, was discovered in the repeated prisoner’s dilemma game [5]. Counterintuitively, ZD strategies unilaterally control payoffs of players by enforcing linear relations between payoffs. ZD strategies can be regarded as generalization of the equalizer strategy [6], which unilaterally sets the payoff of the opponent. ZD strategies also contain the extortionate ZD strategy, by which a player never obtains payoff lower than that of the opponent, and the generous ZD strategy [7], by which a player never obtains payoff higher than that of the opponent but promotes cooperation. So far, ZD strategies were discovered in the prisoner’s dilemma game [5, 7, 8], the public goods game [9, 10], the continuous donation game [11], the asymmetric prisoner’s dilemma game [12, 13], two-player potential games [14, 15], and symmetric games with no generalized rock-paper-scissors cycles [16, 17]. Particularly, ZD strategies in games with infinite action sets are called autocratic strategies [11]. Furthermore, ZD strategies were extended to broader situations, such as games with a discount factor [18, 11, 19, 20], games with observation errors [21, 22, 23], asynchronous games [24], and stochastic games [25, 26]. Roles of ZD strategies in the context of evolution of cooperation have been extensively investigated [27, 28, 7, 29, 30, 31, 8, 32, 33].
@@
Although the Cournot oligopoly game contains some social dilemma structure as noted above, it is not completely the same as social dilemmas. This is because the action sets in the Cournot oligopoly game are infinite sets, whereas the action sets in typical social dilemma games contain only two actions, that is, cooperation and defection. Probably for that reason, ZD strategies have not been discovered in the Cournot oligopoly game, and it is not clear whether players can autocratically control payoffs in the game.
If such unilateral payoff control is possible in the Cournot oligopoly game, a firm can lead the other firms to collusion. In fact, such control is possible by ZD strategies in the repeated prisoner’s dilemma game [5]. If the payoffs of two players are positively correlated in a linear relation enforced by a ZD strategy, increase in payoff of the opponent implies increase in payoff of a player using the ZD strategy. Therefore, if the opponent gradually improves its payoff by learning, the payoff of the player using the ZD strategy also increases up to some local optima. Because ZD strategies do not assume rationality of other players, they are useful even if some players are boundedly rational [34], in contrast to traditional strategies in classical game theory [18].
@@
In this paper, we make two key contributions. First, we prove that ZD strategies (or, autocratic strategies) also exist in the repeated Cournot oligopoly game. Concretely, we show that ZD strategies pinning their own payoffs, positively correlated ZD strategies [35, 36], which are the class containing the extortionate ZD strategy and the generous ZD strategy, and negatively correlated ZD strategies exist in the game. Especially, we find that a fair ZD strategy exists in the game, by which a player unilaterally obtains the average payoff of the opponents. Second, we specify performance of the fair ZD strategy in evolution of collusion. We numerically investigate whether collusion is achieved when the fair ZD strategy is used against adaptively learning players [5, 35, 36]. We find that, in a two-player case, the fair ZD strategy promotes collusion of a learning player, but in a three-player case, it cannot promote collusion.
@@
This paper is organized as follows. In Section 2, we introduce the Cournot oligopoly game and its repeated version. In Section 3, we introduce autocratic strategies, and explain their properties related to the existence. In Section 4, we prove that some types of autocratic strategies exist in the repeated Cournot oligopoly game. In Section 5, we provide numerical results. Section 6 is devoted to discussion. The proofs of all theoretical results are provided in Section 7.
@@ 2 Model
We consider the $N$-player Cournot oligopoly game with $N\geq 2$ [1, 2]. In this game, multiple firms produce the same goods at the same cost. The model assumes that the good’s price is determined by consumer demand, and the price decreases as the total amount of the goods increases. Each firm needs to choose its production while considering production of the other firms. If the production of each firm is too small, it obtains small positive profit. However, if the production of a firm is too large, all firms may obtain negative profit. Therefore, the situation is game-theoretic. In the simplest model, the price linearly decreases as the total amount of the goods increases.
This oligopoly game is mathematically described as follows. The action space of player (firm) $j$ is given by $A_{j}:=[0,x_{\mathrm{max}}]$ with $0<x_{\mathrm{max}}<\infty$ for all $j$. The action of player $j$ is described as $x_{j}\in A_{j}$, which represents production of the goods by player $j$. We collectively write the action profile as $\bm{x}:=\left(x_{1},\cdots,x_{N}\right)$. We also use the notation $\mathcal{A}:=\prod_{j}A_{j}$, $A_{-j}:=\prod_{k\neq j}A_{k}$, and $x_{-j}:=\left(x_{1},\cdots,x_{j-1},x_{j+1},\cdots,x_{N}\right)\in A_{-j}$. When we want to emphasize $x_{j}$ in $\bm{x}$, we write $\bm{x}=(x_{j},x_{-j})$. The payoff of player $j$ when the action profile is $\bm{x}$ corresponds to the profit (the difference between sales and the total production cost) described as
@@
where $a>c>0$, $b>0$, and $\theta(\cdot)$ is the step function. That is, $\left[a-b\sum_{k=1}^{N}x_{k}\right]\theta\left(a-b\sum_{k=1}^{N}x_{k}\right)$ represents a price of the goods, and $c$ represents a cost of production. The price is a non-negative decreasing function of the total production $\sum_{k=1}^{N}x_{k}$, and becomes zero when the total production is too large: $\sum_{k=1}^{N}x_{k}\geq a/b$. In this paper, we assume that $x_{\mathrm{max}}\geq a/b$.
It is known that the Nash equilibrium of this game is $x_{j}=x^{(\mathrm{N})}:=\frac{a-c}{(N+1)b}$ and $s_{j}\left(\bm{x}^{(\mathrm{N})}\right)=\frac{(a-c)^{2}}{(N+1)^{2}b}$ for all $j$, where $\bm{x}^{(\mathrm{N})}:=\left(x^{(\mathrm{N})},\cdots,x^{(\mathrm{N})}\right)$. On the other hand, if all firms collude to monopolize the market, and share the payoffs equally, the realized state is $x_{j}=x^{(\mathrm{C})}:=\frac{a-c}{2Nb}$ and $s_{j}\left(\bm{x}^{(\mathrm{C})}\right)=\frac{(a-c)^{2}}{4Nb}$ for all $j$, where $\bm{x}^{(\mathrm{C})}:=\left(x^{(\mathrm{C})},\cdots,x^{(\mathrm{C})}\right)$. Because $s_{j}\left(\bm{x}^{(\mathrm{C})}\right)>s_{j}\left(\bm{x}^{(\mathrm{N})}\right)$, the game can be regarded as one example of social dilemmas. In addition, the Walras equilibrium, which corresponds to perfect competition, is $x_{j}=x^{(\mathrm{W})}:=\frac{a-c}{Nb}$ and $s_{j}\left(\bm{x}^{(\mathrm{W})}\right)=0$ for all $j$, where $\bm{x}^{(\mathrm{W})}:=\left(x^{(\mathrm{W})},\cdots,x^{(\mathrm{W})}\right)$. In perfect competition, firms cannot influence the price, and the price $\left[a-b\sum_{k=1}^{N}x_{k}\right]\theta\left(a-b\sum_{k=1}^{N}x_{k}\right)$ is equal to the marginal cost $c$. In the limit $N\rightarrow\infty$, the total production in the Nash equilibrium converges to $Nx^{(\mathrm{N})}\rightarrow(a-c)/b$, where $(a-c)/b=Nx^{(\mathrm{W})}$ is the total production in perfect competition. This result implies that perfect competition is realized if the number of firms is very large, and is known as the Cournot limit theorem.
We consider a repeated version of the Cournot oligopoly game. We write the probability measure of an action profile at time $t$ by $P_{t}\left(\cdot\right)$. The payoff of player $j$ in the repeated game is given by
@@
where $\left\langle\cdots\right\rangle^{*}$ is the expected value of the quantity $\cdots$ with respect to the limit probability of an action profile
$P^{*}\left(\cdot\right)$ $:=(1-\delta)\sum_{t=1}^{\infty}\delta^{t-1}P_{t}\left(\cdot\right)$ (3)
and $\delta$ is the discount factor satisfying $0\leq\delta<1$. If the limit $\lim_{T\rightarrow\infty}\sum_{t=1}^{T}P_{t}\left(\cdot\right)/T$ exists, we obtain
$\lim_{\delta\rightarrow 1^{-}}P^{*}\left(\cdot\right)$ $=\lim_{T\rightarrow\infty}\frac{1}{T}\sum_{t=1}^{T}P_{t}\left(\cdot\right).$ (4)
In this paper, we focus only on the undiscounted case $\delta\rightarrow 1^{-}$.
@@ 3 Preliminaries
A memory-one strategy of player $j$ is an autocratic strategy when its strategy $T_{j}$ can be written in the form
$\int\psi(x_{j})dT_{j}\left(x_{j}|\bm{x}^{\prime}\right)-\psi\left(x^{\prime}_{j}\right)$ $=\sum_{k=0}^{N}\alpha_{k}s_{k}\left(\bm{x}^{\prime}\right)$ (5)
with some coefficients $\left\{\alpha_{k}\right\}$ and some bounded function $\psi(\cdot)$.
Surprisingly, the autocratic strategy (5) unilaterally enforces a linear relation between payoffs
$0$ $=\sum_{k=0}^{N}\alpha_{k}\left\langle s_{k}\right\rangle^{*}.$ (6)
Therefore, autocratic strategies can be used to control payoffs unilaterally. The left-hand side of Eq. (5) is the difference between an integral of the function $\psi$ with respect to a memory-one strategy $T_{j}$ and an integral of the function $\psi$ with respect to the Repeat strategy [8]. The general meaning of the function $\psi$ is not clear, and it has been heuristically obtained [16, 15]. For two-player potential games, it is related to potential functions [14]. Recently, for stochastic games, the method for finding $\psi$ numerically was proposed [37]. It is also noteworthy that autocratic strategies are equally powerful in multi-player games although they were originally introduced in two-player games [11], because their proof did not use the number of players. In this paper, we use the two words “ZD strategy” and “autocratic strategy” interchangeably.
@@
$0$ $\leq B\left(\overline{x}_{j},x_{-j}\right)\leq W\quad(\forall x_{-j}\in A_{-j}).$ (7)
@@
$\int_{x_{j}\in A_{j}^{\prime}}\psi(x_{j})dT_{j}\left(x_{j}|\bm{x}^{\prime}\right)-\psi\left(x^{\prime}_{j}\right)$ $=B\left(\bm{x}^{\prime}\right)\quad\left(\forall x^{\prime}_{j}\in A_{j}^{\prime},\forall x^{\prime}_{-j}\in A_{-j}\right)$ (10)
with $\psi\left(\underline{x}_{j}\right)=\psi_{0}+W$ and $\psi\left(\overline{x}_{j}\right)=\psi_{0}$ $(\psi_{0}\in\mathbb{R})$. McAvoy and Hauert called such autocratic strategies two-point autocratic strategies, because such an autocratic strategy uses only two actions. This proposition on two-point autocratic strategies is useful, since we can easily construct an autocratic strategy only by specifying the two actions $\underline{x}_{j}$ and $\overline{x}_{j}$.
Although they found that the condition (7) is a sufficient condition for the existence of general autocratic strategies, Ueda proved that it is also a necessary condition when the number of actions of each player is finite [16]. However, we do not know whether it is also a necessary condition when the number of actions of some player is infinite. At this stage, we only prove the following proposition about two-point autocratic strategies.
@@
If a two-point autocratic strategy (which uses only two actions) of player $j$ controlling $B$ exists, then there exist two actions $\underline{x}_{j},\overline{x}_{j}\in A_{j}$ such that